THE LOW FREQUENCY BROADBAND FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF HEXAFLUOROPROPYLENE OXIDE, CF_{3}CFOCF_{2}

\underline{LU KANG}^{a}, Department of Natural Sciences, Union College, Barbourville, KY 40906; STEVEN T. SHIPMAN, JUSTIN L. NEILL, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904; ALBERTO LESARRI, Departmento de Quimica Fisica, Universidad de Valladolid, Facultad de Ciencias, Valladolid, 47005 SPAIN; and BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904.

The pure rotational spectra of hexafluoropropylene oxide (HFPO), CF_{3}CFOCF_{2}, as well as its ^{13}C (1.07%) and ^{18}O (0.205%) isotopomers were recorded in natural abundance from 2.0 to 26 GHz. Low frequency transitions (2 - 8 GHz) were measured by a recently designed chirped-pulse Fourier transform microwave (CP-FTMW) spectrometer at the University of Virginia. The observed spectra lines of ^{13}C isotopomers (in natural abundance) demonstrate the capability and sensitivity of the CP-FTMW spectrometer operating in 2 - 8 GHz region. The spectra in 8.0 - 26 GHz region were recorded by a Fabry-Pérot cavity Fourier transform microwave (FP-FTMW) spectrometer. All the five isotopologues of HFPO were found, and their spectroscopic constants were fit from assigned spectral lines using JB95 and Pickett^{,}s SPFIT suite of programs. For the dominate HFPO isotopomer: A_{0} = 2217.04887(11) MHz, B_{0} = 1101.48958(5) MHz, C_{0} = 936.60131(5) MHz, Δ_{J} = 55.0(2) Hz, Δ_{JK} = 107.5(9) Hz, Δ_{K} = -20(2) Hz, δ_{J} = 8.49(6) Hz, and δ_{K} = -266(2) Hz. The experimentally determined molecular structure and rotational constants are in a good agreement with our density functional theory calculation using B3LYP/6-31g(d) method.

aThe John B. Stephenson Fellowship of ACA is gratefully acknowledged.